# Homepage: https://kongzhecn.github.io/omg-project/ Code: https://github.com/kongzhecn/OMG/

### v1

@@ Abstract
Image Generation Multi-Concept Customization Diffusion Model
@@ 1 Introduction
Several multi-concept personalization methods have been proposed [25, 41, 28, 15], but they still encounter identity degradation problems when generating multiple concepts. Mix-of-show [16] can generate multi-concepts with realistic identity, but it cannot handle occlusion between concepts. Specifically, the method [16] adopts a regionally controllable sampling method, where each timestep injecting region prompts through regional-aware cross-attention. In cases where the concept regions experience occlusion, the final prediction results for these occluded regions are determined by a straightforward linear addition of the cross-attention results from multiple local sample regions. This simplistic approach leads to inaccurate predictions within the occluded regions, resulting in layout conflicts and identity degradation, as shown in Fig. 2 (b). Besides, there is disharmony between the foreground and background, leading to unnatural illumination in the image. Additionally, methods [16, 25] aims at merging two concepts into one diffusion model, which is computationally inefficient.
@@ 3.1 Preliminary
Latent diffusion model [21, 39, 32] belongs to a class of generative models containing a diffusion process and a reverse process in the latent space. In the diffusion process, an image $x$ is firstly projected to latent space by an encoder $\mathcal{E}$: $z_{0}=\mathcal{E}(x)$. Then random Gaussian noises are gradually added to the data sample $z_{0}$ to generate the noisy sample $z_{t}$ with a predefined noise adding schedule $\alpha_{t}$ as time-step t: $q(z_{t}|z_{0})=\mathcal{N}(\sqrt{\bar{\alpha}_{t}}z_{0},(1-\bar{\alpha}_{t})I)$, where $\bar{\alpha}_{t}=\textstyle\prod_{i=1}^{t}\alpha_{i}$. In the reverse process, a U-Net $\varepsilon_{\theta}$ is trained to directly perform denoising in the latent space. The overall training objective is defined as
@@ 3.3 Stage 2: Multi-concept Personalized Denoising
Concept Noise Blending. To mitigate the additional optimization costs associated with network merging, the proposed concept noise blending method directly leverages multiple single-concept models during inference, circumventing the need for network merging. Moreover, each single-concept model is solely responsible for generating a specific concept, effectively addressing the challenge of identity degradation.
@@
In each timestep, we ensure that the layout is preserved in the generated image by modifying the cross-attention maps within the UNet during the $T2I$ model sampling. For instance, at the $t$ timestep, $z_{t}$ is fed into the $T2I$ model alongside the global prompt $p$ and timestep $t$. Cross-attention maps play a crucial role in controlling the structure and geometry of an image. To maintain an occlusion-friendly layout, we overwrite the generated attention map in each timestep within the UNet with the stored maps. This process can be formulated as:
@@ 3.4 Denoising Timestep of Concept Noise Blending
The initiation timestep for concept noise blending holds significant influence over both the image layout and illumination of the generated image. To elucidate the impact of different concept noise blending starting points, we present generated images at various timesteps in Fig. 5. Leveraging DDIM, the series comprises a total of $50$ steps, with the leftmost image representing the outcome when Concept Noise Blending begins at step $50$, indicating that concept noise blending operations are active throughout the entire sampling process. The rightmost image represents the result starting from step 00, indicating that no concept noise blending operations occur during the entire sampling process. Hence, when the concept noise blending operation starts at timestep 00, the generated image is the same as stage one. The intermediary images illustrate the progressive steps from $50$ to 00.
Commencing concept noise blending at an early step may introduce layout conflicts in the composition and shape of objects within the generated image. However, with the increase in concept noise blending steps, the content information becomes more coherent and stable, effectively preserving the identity of the object. After iterative denoising, as the concept noise blending step approaches 00, the identity of the character diminishes gradually, resulting in a synthesized image resembling the first stage. This highlights the early stage of sampling governs the image layout, while the identity of concepts unfolds in later timesteps.
@@ 4 Experiments
Figure 7: Comparison with InstantID [44] in single-concept customization. OMG emerges as the superior method by generating images with more natural colors. This showcases the prowess of OMG over InstantID in the context of single-concept customization.
@@ 4.1 Quantitative Comparison
Following Custom Diffusion [25], we utilize $20$ text prompts and $50$ samples per prompt for each concept. Hence, a total of $1000$ images are ultimately generated. For a fair comparison, all the comparison methods adopt DDPM sampling with $50$ steps and a classifier-free guidance sample across all methods. Our evaluation spans various categories of concepts, including characters and objects. We use a single-concept tuned model to assess the identity-preserving effect of our method through a set of prompts. The experimental results including single-concept and multi-concepts are detailed in Tab. 1.
@@
Table 2: Quantitative comparison on multi-concept personalization. OMG exhibits favorable generative effects on both characters simultaneously
Method Identity Alignment
Character 1 Character 2 Average
DreamBooth [33] 0.136 0.493 0.315
Textual Inversion [12] 0.023 0.228 0.126
Custom Diffusion [25] 0.098 0.252 0.175
Mix-of-show [16] 0.357 0.257 0.307
OMG (Ours) 0.382 0.478 0.430
The qualitative results for multi-concept personalization shown in Tab. 1 mainly measure the fusion ability of multiple single-concept models during multi-concept customization. It cannot reflect the generation effect of multi-concept generation. Therefore, to measure the generation effects when different methods generate multiple concepts simultaneously, we propose a new calculation method. To make a fair comparison, we use the same spatial condition to generate two characters simultaneously, one male and another female. We calculate the region of different characters in the image through visual comprehension, then calculate the Identity Alignment scores for two different characters with their corresponding reference images separately. This measure approach is more effective in measuring the effects of generating multiple concepts simultaneously. The experiment results are shown in Tab. 2. Some comparison methods may derive good Identity Alignment in one character but not perform well on another. Our method achieves satisfied generative effects on both characters simultaneously and obtains the best Average Identity Alignment. This demonstrates the effectiveness of our method in multi-concept generation.

### v2

@@ Abstract
Image Generation Image Customization Diffusion Model
@@ 1 Introduction
Several multi-concept personalization methods have been proposed [25, 41, 28, 15], but they still encounter identity degradation problems when generating multiple concepts. Mix-of-show [16] can generate multi-concepts with realistic identity, but it cannot handle occlusion between concepts. Specifically, the method [16] adopts a regionally controllable sampling method, where each timestep injecting region prompts through regional-aware cross-attention. In cases where the concept regions experience occlusion, the final prediction results for these occluded regions are determined by a straightforward linear addition of the cross-attention results from multiple local sample regions. This simplistic approach leads to inaccurate predictions within the occluded regions, resulting in layout conflicts and identity degradation, as shown in Fig. 2 (b). Besides, there is disharmony between the foreground and background, leading to unnatural illumination in the image. Additionally, methods [16, 25] aim at merging two concepts into one diffusion model, which is computationally inefficient.
@@ 3.1 Preliminary
Latent diffusion model [21, 39, 32] belongs to a class of generative models containing a diffusion process and a reverse process in the latent space. In the diffusion process, an image $x$ is firstly projected to latent space by an encoder $\mathcal{E}$: $z_{0}=\mathcal{E}(x)$. Then random Gaussian noises are gradually added to the data sample $z_{0}$ to generate the noisy sample $z_{t}$ with a predefined noise adding schedule $\alpha_{t}$ as timestep t: $q(z_{t}|z_{0})=\mathcal{N}(\sqrt{\bar{\alpha}_{t}}z_{0},(1-\bar{\alpha}_{t})I)$, where $\bar{\alpha}_{t}=\textstyle\prod_{i=1}^{t}\alpha_{i}$. In the reverse process, a U-Net $\varepsilon_{\theta}$ is trained to directly perform denoising in the latent space. The overall training objective is defined as
@@ 3.3 Stage 2: Multi-concept Personalized Denoising
Concept Noise Blending. To mitigate the additional optimization costs associated with network merging, the proposed method directly leverages multiple single-concept models during inference, circumventing the need for network merging. Moreover, each single-concept model is solely responsible for generating a specific concept, effectively addressing the challenge of identity degradation.
@@
In each timestep, we ensure that the layout is preserved in the generated image by modifying the cross-attention maps within the UNet during the $T2I$ model sampling. For instance, at the $t$ step, $z_{t}$ is fed into the $T2I$ model alongside the global prompt $p$ and timestep $t$. Cross-attention maps play a crucial role in controlling the structure and geometry of an image. To maintain an occlusion-friendly layout, we overwrite the generated attention map in each timestep within the UNet with the stored maps. This process can be formulated as:
@@ 3.4 Denoising Timestep of Concept Noise Blending
The initiation timestep for concept noise blending holds significant influence over both the image layout and illumination of the generated image. To elucidate the impact of different concept noise blending starting points, we present generated images at various timesteps in Fig. 5. Leveraging DDIM, the series comprises a total of $50$ steps, with the leftmost image representing the outcome when Concept Noise Blending begins at step $50$, indicating that concept noise blending operations are active throughout the entire sampling process. The rightmost image represents the result starting from step 00, indicating that no concept noise blending operations occur during the entire sampling process. Hence, when the concept noise blending operation starts at timestep 00, the generated image is the same as stage one.
Commencing concept noise blending at an early step may introduce layout conflicts in the composition and shape of objects within the generated image. However, with the increase in concept noise blending steps, the content information becomes more coherent and stable, effectively preserving the identity of the object. After iterative denoising, as the concept noise blending step approaches 00, the identity of the character diminishes gradually, resulting in a synthesized image resembling the first stage. This highlights the early stage of sampling governs the image layout, while the identity of concepts unfolds in later timesteps. The optimal step of concept noise blending is approximately $35$.
@@ 4 Experiments
Figure 7: Comparison with InstantID [44] in single-concept customization. OMG emerges as the superior method by generating images with more natural colors. This showcases the prowess of OMG over InstantID in single-concept customization. For LoRA and InstantID, we all adopt SDXL-base-1.0 as the base model for a fair comparison.
@@ 4.1 Quantitative Comparison
Following Custom Diffusion [25], we utilize $20$ text prompts and $50$ samples per prompt for each concept. Hence, a total of $1000$ images are ultimately generated. For a fair comparison, all the comparison methods adopt DDIM sampling with $50$ steps and a classifier-free guidance sample across all methods. Our evaluation spans various categories of concepts, including characters and objects. We use a single-concept tuned model to assess the identity-preserving effect of our method through a set of prompts. The experimental results including single-concept and multi-concepts are detailed in Tab. 1.
@@
Table 2: Quantitative results for diverse concepts generation.
Method man + woman man + man woman + woman object + object Average
IDA IDA IDA IMA
DreamBooth [33] 0.302 0.258 0.192 0.784 0.384
Textual Inversion [12] 0.122 0.131 0.064 0.675 0.248
Custom Diffusion [25] 0.265 0.210 0.212 0.757 0.361
Mix-of-show [16] 0.361 0.219 0.143 0.736 0.365
OMG (Ours) 0.487 0.377 0.293 0.763 0.480
The qualitative results for multi-concept personalization shown in Tab. 1 mainly measure the fusion ability of multiple single-concept models during multi-concept customization. It cannot reflect the generation effect of multi-concept generation. Therefore, to measure the generation effects when different methods generate multiple concepts simultaneously, we propose a new calculation method. To make a fair comparison, we use the same spatial condition to generate two concepts simultaneously. We calculate the region of different concepts in the image through visual comprehension, then calculate the average Identity Alignment scores (IDA) or Image Alignment score (IMA) for two different concepts with their corresponding reference images separately. This measure approach is more effective in measuring the effects of generating multiple concepts simultaneously. The experiment results are shown in Tab. 2. OMG outperforms other methods in generating images with various concept combinations, demonstrating the effectiveness of our method in multi-concept generation.
@@ 5 Conclusion
Acknowledgment
This work is funded in part by the National Natural Science Foundation of China (Grant No. 62372480), in part by CCF-Tencent Rhino-Bird Open Research Fund (No. CCF-Tencent RAGR20230118), in part by Theme-based Research Scheme (T45-205/21-N) from Hong Kong RGC, in part by Generative AI Research and Development Centre from InnoHK.